OpenStudy (anonymous):

let f(x) = 3x^2 +8x-16 and g(x)=x+4. perform the function operation and then find the domain of the result. (f . g) (x) (this is a compostition)

OpenStudy (anonymous):

Put g(x) in for every x in f(x). \[f(g(x))=3(x+4)^{2}+8(x+4)-16\] f(g(x))=3x^2 + 32x + 64 The domain is \[-\infty < x < \infty\]

OpenStudy (anonymous):

so is the domain all real numbers? or is it either x < 0 or x >0

OpenStudy (anonymous):

All real numbers (:

OpenStudy (anonymous):

thank you!!!!